You don't want to spend the winter shovelling snow so you decide to hire a local snow removal company. You know that most companies charge an initial fee and a fixed amount each time that they come to clear your driveway.
Olivia recommends that you use Company A. She doesn't remember the initial fee, but knows that it cost $20 Every time they came,she paid a total of $390 by the end of the winter and they came 17 times.
Kevin recommends that you use Company B. There was an initial fee of $125 but then it only cost $15 each time they came to clear the driveway.
When will it cost the same amount to choose either company? Which company would you choose and Why?

Answer by Theo(13342)

she paid 20 dollars each time they came and they came 17 times.
the cost was 20 * 17 = 340.
she paid 390, so it looks like company A had an initial fee of 50 dollars and charged an additional 20 dollars for each visit.

company B had an initial fee of 125 dollars and charged an additional 15 dollars for each visit.

the cost will be the same when 50 + 20 * x = 125 + 15 * x
x represents the number of visits.
subtract 50 from both sides of the equation and subtract 15 * x from both sides of the equation to get:
20 * x - 15 * x = 125 - 50
simplify to get:
5 * x = 75
solve for x to get:
x = 15
the cost will be the same when 15 visits have been made.
company A will charge 125 + 15 * 15 = 350 dollars.
company B will charge 50 + 15 * 20 = 350 dollars.

if she expects less than 15 visits, then she should go with company A.
if she expects more than 15 visits, then she should go with company B.

for example:
at 10 visits:
company B would charge 125 + 10 * 15 = 275
company A would charge 50 + 10 * 20 = 250
company A would be cheaper.

at 20 visits:
company B would charge 125 + 20 * 15 = 425
company A would charge 50 + 20 * 20 = 450
company B would be cheaper.

they equations can be graphed.
in that way you can visualize where the break even point is and which company is better before the breaking point and which company is better after the breaking point.